Given F(X)=x-3 and g(X)= X^2 over x-3 find g(f(x))
Answer by nyc_function(2741)

g(f(x)) = plug value of f(x) into the value of g(x) and simplify.
We know the value of f(x) is x - 3.
Everywhere you see x for the function g(x), replace it with x - 3 and simplify.
g(x - 3) = (x - 3)^2/(x - 3) - 3
g(x - 3) = (x - 3)(x - 3)/(x - 6)
So, g(f(x)) = (x - 3)(x - 3)/(x - 6)